OpenStudy (anonymous):

Would love some help on a few of these questions What is the simplest form of sqrt{80x^7y^6} What is the simplest form of sqrt[3]{25xy^2}*sqrt[3]{15x^2}

OpenStudy (anonymous):

\[\sqrt{80x^7y^6}\] \[\sqrt[3]{25xy^2}*\sqrt[3]{15x^2}\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Sorry, I'm horrible at math!

OpenStudy (anonymous):

Oh that's ok I just saw you replying to another question ^.^

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

@Michele_Laino @Catlover5925

OpenStudy (catlover5925):

@sammixboo will help she is awesome

OpenStudy (michele_laino):

we can write: \[\sqrt{80x ^{7}y ^{6}}=\sqrt{16*5*x ^{6}*y ^{6}*x}=4x ^{3}y ^{3}\sqrt{5x}\]

sammixboo (sammixboo):

I don't know for #1 but for #2 I know the answer, but I can't explain it =.= Gimme a second

OpenStudy (anonymous):

these are my answer/questions

OpenStudy (michele_laino):

second question: \[\sqrt[3]{25xy ^{2}}*\sqrt[3]{15x ^{2}}=\sqrt[3]{25*15x ^{3}y ^{2}}=\] \[\sqrt[3]{125*3*x ^{3}y ^{2}}=\sqrt[3]{5^{3}x ^{3}y ^{2}}=5x \sqrt[3]{y ^{2}}\]

OpenStudy (anonymous):

Fantastic! Thank you @Michele_Laino do you have time for a few more?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Fantastic! I'l make a new thingy